OpenStudy (anonymous):

What is the solution to the equation 9^(x – 1) = 27 ?

OpenStudy (anonymous):

it cant be 4, i already checked

OpenStudy (anonymous):

x = 2.5 x = 0.5 x = –0.5 x = –3.5 has to be one of these 4 choices

OpenStudy (anonymous):

it is

OpenStudy (campbell_st):

rewrite it \[(3^2)^{x - 1} = 3^3\] looking at the powers 2x - 2 = 3 2x = 5 x = 5/2

OpenStudy (anonymous):

i got it now, thanks guys!